OpenStudy (anonymous):

Write a rule for the linear function in the table. To write a function rule fro a table look for a pattern. The slope m is f(x)/x, and b is the value of f(x) when x=0. Help please I'm confused. x f(x) -4 16 0 0 4 -16 8 -32

OpenStudy (precal):

use the slope formula and two points y=mx+b because of (0,0) your b value is 0 you just need your m value

OpenStudy (anonymous):

um would that be m= 16/4=4 ?

OpenStudy (precal):

which two points did you use?

OpenStudy (anonymous):

i didn't use two points. what i did was from the table the pattern was ( x side ) 4 and f(x) 16. so my m value = 16/4 = 4. Am I doing this wrong ?

OpenStudy (anonymous):

I'm following this.

OpenStudy (anonymous):

the slope, m, equals the difference in f(x) over the difference in X, NOT f(x) over X. It says this at the top of your screenshot. Note: f(x) is usually stated as Y. Take two points: \[(X_{1},f(x)_{1}) \text{ and } (X_{2},f(x)_{2})\] \[\frac{f(x)_{2}-f(x)_{1}}{X_{2}-X_{1}}=m\] Put the numbers in, solve, and you'll get m.

OpenStudy (anonymous):

thank you

OpenStudy (precal):